OpenStudy (anonymous):

x = -4y + 4 2x + 8y = 8 both put into standard form please

OpenStudy (callisto):

standard form of a linear equation is ax+by+c=0 therefore x = -4y + 4 x+4y-4 = 0 2x + 8y = 8 2x + 8y -8=0

OpenStudy (anonymous):

standard fprm

OpenStudy (anonymous):

simplest form

OpenStudy (callisto):

2x + 8y = 8 2x + 8y -8=0 x+4y-4=0

OpenStudy (anonymous):

not that way the y by itsself

OpenStudy (anonymous):

???????????